The value of a mirror exceeding 10 m2 in area varies directly as the square of its area. Given that a 60 m2 mirror costs $400. Find the price of 45 m2 mirror?
Solution with explanation please

Answer by TimothyLamb(4379)

v = M(a^2)
v value of mirror
a = mirror area
M constant of proportionality for the value of this type of mirror
---
v = M(a^2)
400 = M(60^2)
M = 400/(60^2)
M = 0.111111
---
v = M(a^2)
v = M(45^2)
v = 0.111111(2025)
---
Answer:
224.999999
the value of a 45 sq.m mirror is $225.00
---
the problem statement tells us that the value of this type of mirror varies DIRECTLY (i.e. in the same proportion), as the area-squared of the mirror.
---
the constant of proportionality (M) is necessary in order to properly relate area-squared to value.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php